Perform the indicated operation; write answer in simplest form:
7 and 2/3 – 3 and 1/5
and
2) Perform the indicated operations; write answer in simplest form; note: "•" denotes multiplication:
4 and 2/3 • 3 and 1/3 • 3 and 3/4

Answer by checkley77(12844)

7 and 2/3 – 3 and 1/5
(7*3+2)/3-(3*5+1)/5
(21+2)/3-(15+1)/5
23/3-16/5
(23*5-16*3)/3*5
(115-48)/15
67/15 ans.
2) Perform the indicated operations; write answer in simplest form; note: "•" denotes multiplication:
4 and 2/3 • 3 and 1/3 • 3 and 3/4
(4*3+2)/3*(3*3+1)/3*(3*4+3)/4
(12+2)/3*(9+1)/3*(12+3)/4
14/3*10/3*15/4
14*10*15/3*3*4
2100/36=700/12=350/6=175/3 ans.